given 95 weight% of HCL and Given density is 1.08g/mol

find the molarity of HCL


1
Expert's answer
2021-09-06T02:38:09-0400


•If 1ml contains 1.08g HCL,

then 1000ml (1L) contains (1.08g x1000) = 1080g HCL (assuming the acid is 100% pure)

•If 100% HCL solution contains 1180g, therefore


95% HCL solution contains"= \frac{(95 ×1080)}{(100)} = 1026g/L"


But 36.5g/L of HCL = 1M,

therefore 1026g/L "= \frac{(1026g/L×1M)}{(36.5g/L)} = 28.11M"